What is the chemical formula of Magnesium Ammonium Phosphate and its molar mass?
Answer #1

The chemical formula of anhydrous magnesium ammonium phosphate is (NH4)MgPO4 and molar mass is 137

It can be calculated as = (14.0067*1) + (1.00784*4) + 24.305 + 30.9738+ (15.999*4) = 137.32 g/mol

Mostly it exists as hexahydrate form i.e. (NH4)MgPO4. 6H2O in which case molar mass is 245.41 g/mol (= 137.32 +(18*3) )